# An In-depth Evaluation of GPT-4 in Sentence Simplification with Error-based Human Assessment

### v1

@@ 7.2.2. Distinction between high-quality and low-quality simplifications overall
We examined on all $3590$ simplification outputs in Task $2$. Each simplification output was classified as high quality (labeled as $1$) if it received a high rating (a score of $3$) from at least two out of three annotators across all dimensions (fluency, simplicity, and meaning preservation), and low quality (labeled as 00) otherwise. We computed the correlation coefficients between these classifications and the metric scores. As we discussed in Section 6.2, Newsela is different from other corpora for allowing significant meaning loss to prioritize simplicity. To reduce bias stemming from this particular dataset, we also calculated the correlation after excluding Newsela, i.e., only on Turk and ASSET (denoted as ‘T&A’). For each evaluation, we further divide simplification outputs based on the model (GPT-4 vs. Control-T5) to determine if there are differences in metrics’ capabilities.
Results are summarized in Table 9. For Control-T5 simplification outputs, LENS exhibits some capability to distinguish quality, though it shows limited effectiveness overall. BERTScores indicate a stronger capability to distinguish quality, with correlations ranging from $0.64$ to $0.71$, which notably decrease to $0.33$ to $0.36$ when Newsela-derived simplification outputs are removed. For GPT-4 simplification outputs, neither BERTScores nor LENS are effective at distinguishing quality in simplifications.
These results suggest that while these metrics can detect significant quality variations, they fall short when overall quality is high. Therefore, they may not be suitable or sensitive enough to evaluate advanced LLMs like GPT-4. To facilitate a more intuitive understanding, we also incorporate visualization (see Figure 4). For GPT-4, regardless of the evaluation metric used, the scores of high and low-quality simplification outputs appear to blend, indicating the lack of discriminative capability. In contrast, Control-T5 shows a clear difference, with both LENS and BERTScores tending to give lower scores to low-quality simplification outputs and higher scores to high-quality ones. This differentiation is particularly pronounced in BERTScores, which show a more distinct separation between the evaluation scores of high and low-quality simplification outputs. Note that the majority of these low-quality simplification outputs stem from Newsela, thus this distinction becomes less apparent when excluding Newsela-derived simplifications.
@@
All GPT-4 Control-T5
Raw DS (1515) T&A Raw DS (271) T&A Raw DS (551) T&A
LENS $0.12$ $0.11$ $0.02$ $0.08$ $0.11$ $0.08$ $0.38$ $0.42$ $0.18$
BERT precision $0.34$ $0.35$ $0.23$ $0.10$ $0.19$ $0.13$ $0.67$ $0.71$ $0.33$
BERT recall $0.28$ $0.29$ $0.21$ $0.09$ $0.19$ $0.03$ $0.64$ $0.67$ $0.36$
BERT f1 $0.34$ $0.35$ $0.23$ $0.10$ $0.20$ $0.09$ $0.68$ $0.71$ $0.36$
@@ 7.2.3. Distinction between high-quality and low-quality simplifications within a specific dimension
We examined on $3590$ model-generated simplification outputs in Task $2$ across individual dimensions. For each dimension, simplification outputs were classified as high quality (labeled as $1$) if they received a high rating (a score of $3$) from at least two out of three annotators, and low quality (labeled as 00) otherwise. Based on our classification, on fluency, all the GPT4-generated simplification outputs are high quality, and only five out of $1795$ Control-T5-generated simplification outputs are low quality. Given that GPT-4 and Control-T5 are rare to generate disfluent outputs, we focus on the dimensions of meaning preservation and simplicity. We then computed the correlation between these ratings and metrics scores. We also report results excluding Newsela as in Section 7.2.2.
Table 10. Point-biserial correlation between sentence-level metrics scores and human ratings in single dimension
@@
All GPT-4 Control-T5
Raw DS (1375) T&A Raw DS (145) T&A Raw DS (565) T&A
LENS $0.08$ $0.08$ $-0.07$ $-0.02$ $-0.09$ $-0.03$ $0.37$ $0.40$ $0.16$
BERT precision $0.37$ $0.40$ $0.19$ $0.17$ $0.33$ $0.06$ $0.69$ $0.72$ $0.31$
BERT recall $0.30$ $0.33$ $0.23$ $0.16$ $0.28$ $0.09$ $0.65$ $0.68$ $0.36$
BERT f1 $0.36$ $0.39$ $0.22$ $0.17$ $0.33$ $0.07$ $0.70$ $0.73$ $0.35$
@@
All GPT-4 Control-T5
Raw DS (157) T&A Raw DS (130) T&A Raw DS (27) T&A
LENS $0.11$ $0.27$ $0.16$ $0.14$ $0.32$ $0.10$ $0.04$ $0.18$ $0.16$
BERT precision $-0.05$ $-0.09$ $0.14$ $-0.04$ $-0.06$ $0.12$ $-0.11$ $-0.52$ $0.19$
BERT recall $-0.03$ $-0.06$ $0.05$ $-0.03$ $-0.08$ $0.00$ $-0.07$ $-0.31$ $0.13$
BERT f1 $-0.05$ $-0.08$ $0.11$ $-0.04$ $-0.06$ $0.07$ $-0.10$ $-0.46$ $0.18$
Table 10(a) indicates results for meaning preservation. Overall, BERTScores demonstrate a stronger correlation with human evaluations for meaning preservation compared to LENS scores. Upon dividing the data by model, BertScores show high correlations between $0.65$ and $0.73$ for Control-T5 simplification outputs. However, these correlations significantly drop to between $0.31$ and $0.36$ upon the exclusion of simplification outputs derived from Newsela. For GPT-4 simplification outputs, both metrics reveal a negligible correlation. Task $2$’s human evaluation suggests that Control-T5 significantly underperforms in preserving meaning within the Newsela dataset compared to other datasets. While both metrics effectively differentiate substantial differences in meaning preservation, they struggle when overall meaning preservation is high. Table 10(b) shows results for simplicity, where neither metric correlates well with human evaluations, though LENS slightly outperforms BERTScores. Consequently, they may not be sufficiently sensitive to evaluate advanced LLMs like GPT-4 on meaning preservation and simplicity.
@@ 7.3. Corpus-level Results
Our human evaluations reveal GPT-4’s simplification outputs are generally superior, evidenced by fewer errors, better meaning preservation, and comparable fluency and simplicity to those generated by Control-T5. In this section, we compare the corpus-level metircs scores of the Control-T5 and GPT-4 with human evaluation results to determine if they align, i.e., whether they rate GPT-4 higher than Control-T5.
Table 11. Corpus-level metrics scores for different models
@@
Turk Control-T5 $\mathbf{43.7}$ $68.2$ $*\mathbf{5.8*}$
GPT-4 $42.9$ $*\mathbf{71.5*}$ $8.3$
ASSET Control-T5 $44.9$ $*\mathbf{74.5*}$ $*\mathbf{6.3*}$
GPT-4 $*\mathbf{47.3*}$ $59.3$ $7.6$
Newsela Control-T5 $38.6$ $*\mathbf{23.5*}$ $*\mathbf{4.2*}$
GPT-4 $*\mathbf{41.4*}$ $13.7$ $5.7$
@@
The metrics’ scores are detailed in Table 11, with the better scores emphasized in bold. To assess the statistical significance of the differences in corpus-level scores, we employed a randomization test (Fisher, 1935) against Control-T5. Better corpus-level scores are masked with an asterisk (***) if statistically significant differences were confirmed. SARI favors GPT-4 over Control-T5 in ASSET and Newsela, aligning with our human evaluations of overall superior performance. In Turk, both models are scored similarly. Although overall GPT-4 surpasses Control-T5 in Turk, our error analysis points out that a significant portion of the errors in GPT-4’s outputs stem from the use of more complex lexical expressions. SARI appears sensitive on lexical simplicity, the same as what was reported in a previous study (Xu et al., 2016). BLEU significantly favors Control-T5 on ASSET and Newsela, which does not match our human evaluations. Studies (Xu et al., 2016; Sulem et al., 2018a) have demonstrated that BLEU is unsuitable for simplification tasks, as it tends to negatively correlate with simplicity, often penalizing simpler sentences, and gives high scores to sentences that are close or even identical to the input. Our findings further underscore the limitations of BLEU in evaluating sentence simplification. FKGL ranks Control-T5’s outputs as easier to read compared to those from GPT-4 across all datasets. This aligns with our human evaluation; Control-T5 tends to generate simpler sentence expense of meaning preservation, thereby making the sentence easier to read. However, FKGL’s focus solely on readability, without taking into account the quality of the content or the reference sentences, limits its effectiveness in a comprehensive quality analysis. Previous studies (Alva-Manchego et al., 2021; Tanprasert and Kauchak, 2021) show FKGL is unsuitable for sentence simplification evaluation. Our findings further highlight its limitations in accurately evaluating corpus-level sentence simplification.
@@ 7.4. Summary of Findings
We summarize our findings on the meta-evaluation of existing evaluation metrics for sentence simplification, namely LENS, BERTScores, SARI, BLEU, and FKGL.
@@
    At sentence level, BERTScores are effective in differentiating the distinct difference between high-quality and low-quality simplification overall as well as on meaning preservation. However, they are not sensitive enough to evaluate the quality of simplification generated by GPT-4.
@@
    On sentence-level simplicity, neither of the metrics is capable of differentiating high-quality and low-quality simplification.
@@
    At corpus level, SARI aligns with our human evaluations while BLEU does not.
@@ 8. Conclusion
In this study, we conducted an in-depth human evaluation of the advanced LLM, specifically the performance of GPT-4, in sentence simplification. Our findings highlight that GPT-4 surpasses the supervised baseline model, Control-T5, by generating fewer erroneous simplification outputs and preserving the source sentence’s meaning better. These results underscore the superiority of advanced LLMs in this task. Nevertheless, we observed limitations, notably in GPT-4’s handling of lexical paraphrasing. Further, our meta-evaluation of sentence simplification’s automatic metrics demonstrated their inadequacy in accurately assessing the quality of GPT-4-generated simplifications.
Our investigation opens up multiple directions for future research. Future studies could investigate how to mitigate lexical paraphrasing issues. For example, it would be worthwhile to explore whether fine-tuning could help. Moreover, there’s a need for more sensitive automatic metrics to properly evaluate the sentence-level quality of simplifications generated by LLMs.
@@ A.1. Best Prompts in GPT-4’s prompt engineering
Figure 5 illustrates the best prompts that achieved the highest SARI scores on each validation set during GPT-4’s prompt engineering. Each prompt comprises: instructions, examples of original to simplification(s) transformation, and a source sentence.

### v2

@@ 7.2.2. Distinction Between High-Quality and Low-Quality Simplifications Overall
We examine all $10,471$ model-generated simplification outputs in Task $2$. Each simplification output is classified as high quality (labeled as $1$) if it received a high rating (a score of $3$) from at least two out of three annotators across all dimensions (fluency, simplicity, and meaning preservation), and low quality (labeled as 00) otherwise. We compute the correlation coefficients between these classifications and the metric scores. As we discussed in Section 6.2, Newsela is different from other corpora in that it allows significant meaning loss to prioritize simplicity. To reduce bias stemming from this particular dataset, we also calculate the correlation after excluding Newsela, i.e., only on Turk, ASSET, and /or SimPA (denoted as ‘T&A&S’ for overall and LLMs, and ‘T&A’ for Control-T5). The distributions of high-quality and low-quality outputs are generally highly imbalanced. For example, GPT-4 generates $2,593$ overall high-quality simplification outputs compared to just $299$ low-quality ones, potentially affecting correlation results. To address this, we only report correlations after downsampling. For each evaluation, we further divide simplification outputs based on the model to determine if there are differences in the metrics’ capabilities. The results are summarized in Table 6.
Metrics Fail to Effectively Differentiate Between High- and Low-Quality
LENS shows some ability to distinguish quality in Control-T5 simplification outputs ($0.34$ of correlation coefficient) but remains limited overall ($0.15$ in ‘All-DS’, and 0.010.010.010.010.340.340.340.34 across models). BERTScore generally performs better across all models, though its effectiveness is not reliable enough ($0.34$ in ‘All-DS’, and 0.200.200.200.200.600.600.600.60 across models). Both metrics exhibit slightly higher correlation when evaluating Control-T5 (LENS: $0.34$, BERT f1: $0.60$) and Llama-3.2-3B (LENS: $0.15$, BERT f1: $0.39$) compared to GPT-4 (LENS: $0.01$, BERT f1: $0.20$) and Qwen2.5-72B (LENS: $0.03$, BERT f1: $0.33$). However, their performance declines notably after removing Newsela-derived simplification outputs.
More Challenges in Evaluating High-quality Simplification Models
To further compare the metrics’ evaluations of high- and low-quality outputs across models, we incorporate visualizations (see Figure 6) after downsampling. For GPT-4 and Qwen2.5-72B, regardless of the evaluation metric used, the scores of high and low-quality simplification outputs appear to blend, revealing a lack of discriminative capability. This indicates that these metrics struggle to differentiate when the overall quality is high, making them less suitable for evaluating advanced LLMs like GPT-4 and Qwen2.5-72B. In contrast, for Llama-3.2-3B and Control-T5, high-quality sentence pairs rated by humans tend to receive higher scores from both metrics, with LENS showing a clearer alignment. However, low-quality sentence pairs rated by humans exhibit a wider range of scores, showing the metrics’ limitations in capturing quality variations.
@@
All GPT-4 Qwen2.5-70b Llama-3.2-3B Control-T5
DS (3,164) T&A&S-DS (936) DS (299) T&A&S-DS (104) DS (309) T&A&S-DS (72) DS (1,312) T&A&S-DS (530) DS (551) T&A-DS (230)
LENS 0.15 0.01 0.01 0.06 0.03 0.18 0.15 -0.07 0.34 0.18
BERT f1 0.34 0.16 0.20 -0.07 0.33 0.21 0.39 0.27 0.60 0.37
@@ 7.2.3. Distinction Between High-Quality and Low-Quality Simplifications Within a Specific Dimension
We examined $10,471$ model-generated simplification outputs in Task $2$ across individual dimensions. For each dimension, simplification outputs are classified as high quality (labeled as $1$) if they received a high rating (a score of $3$) from at least two out of three annotators, and low quality (labeled as 00) otherwise. Based on our classification, on fluency, only one GPT4, one Qwen2.5-70b, two Llama-3.2-3B, and five Control-T5-generated simplification outputs are low quality. Given that these models rarely generate disfluent outputs, we focus on the dimensions of meaning preservation and simplicity. We then compute the correlation between these ratings and metrics scores. As in Section 7.2.2, we only report results after downsampling and incorporate results after excluding Newsela.
Table 7. Point-biserial correlation between sentence-level scores and human ratings in single dimension
@@
All GPT-4 Qwen2.5-72B Llama-3.2-3B Control-T5
DS ($2,848$) T&A&S-DS ($767$) DS ($168$) T&A&S-DS ($37$) DS ($263$) T&A&S-DS ($43$) DS ($1,187$) T&A&S-DS ($471$) DS ($565$) T&A-DS ($216$)
LENS $0.11$ $-0.09$ $-0.04$ $-0.02$ $-0.02$ $0.10$ $0.06$ $-0.06$ $0.31$ $0.19$
BERT f1 $0.37$ $0.24$ $0.40$ $0.30$ $0.48$ $0.33$ $0.42$ $0.34$ $0.58$ $0.38$
@@
All GPT-4 Qwen2.5-72B Llama-3.2-3B Control-T5
DS ($451$) T&A&S-DS ($188$) DS ($135$) T&A&S-DS ($68$) DS ($47$) T&A&S-DS ($30$) DS ($242$) T&A&S-DS ($65$) DS ($27$) T&A-DS ($25$)
LENS $0.43$ $0.28$ $0.19$ $0.21$ $0.43$ $0.48$ $0.57$ $0.23$ $0.16$ $0.31$
BERT f1 $0.06$ $-0.29$ $-0.04$ $-0.33$ $-0.02$ $-0.06$ $0.26$ $-0.39$ $-0.19$ $0.50$
Table 7(a) indicates results for meaning preservation. Overall, while moderate, BERTScore shows a stronger correlation with human evaluations of meaning preservation ($0.37$ in ‘All-DS’ and 0.400.400.400.400.580.580.580.58 across models) compared to LENS scores ($0.11$ in ‘All-DS’ and 0.040.04-0.04- 0.040.310.310.310.31 across models). However, these correlations decline from 0.370.370.370.370.580.580.580.58 to $0.24$–-$0.38$ when simplification outputs from Newsela are excluded. Table 7(b) shows results for simplicity. Overall, LENS demonstrates a stronger correlation with human evaluations for simplicity ($0.43$ in ‘All-DS’, and 0.160.160.160.160.570.570.570.57 across models) compared to BERT score ($0.06$ in ‘All-DS’, and 0.190.19-0.19- 0.190.260.260.260.26 across models ).
@@ 7.3. Corpus-level Results
Our human evaluations reveal GPT-4 and Qwen2.5-72B’s simplification outputs are generally superior, evidenced by fewer errors, better meaning preservation, and comparable fluency and simplicity to those generated by Llama-3.2-3B and Control-T5. In this section, we compare the corpus-level metrics scores of the models with human evaluation results to determine if they align, i.e., whether they rate GPT-4 and Qwen2.5-72B higher than Llama-3.2-3B and Control-T5.
Table 8. Corpus-level scores for different models
@@
Turk GPT-4 $42.9$ $\mathbf{72.0}$ $8.3$
Qwen2.5-72B $42.7$ $63.6$ $7.8$
Llama-3.2-3B $38.2$ $55.8$ $7.5$
Control-T5 $\mathbf{43.7}$ $68.2$ $\mathbf{5.8}$
ASSET GPT-4 $47.3$ $59.3$ $7.6$
Qwen2.5-72B $\mathbf{47.9}$ $69.8$ $8.2$
Llama-3.2-3B $45.5$ $67.8$ $8.2$
Control-T5 $44.9$ $\mathbf{74.5}$ $\mathbf{6.3}$
Newsela GPT-4 $41.4$ $13.9$ $5.7$
Qwen2.5-72B $41.7$ $17.0$ $6.3$
Llama-3.2-3B $\mathbf{41.9}$ $13.9$ $\mathbf{4.1}$
Control-T5 $38.6$ $\mathbf{24.0}$ $4.2$
SimPA GPT-4 $40.8$ $23.5$ $9.6$
Qwen2.5-72B $\mathbf{42.8}$ $\mathbf{28.0}$ $10.3$
Llama-3.2-3B $38.2$ $18.9$ $\mathbf{8.5}$
Table 9. Comparison of sentence simplification models and LLMs on the Turk corpus. The first section presents SARI scores measured by Alva-Manchego et al. (Alva-Manchego et al., 2020b), while the second section shows SARI scores for LLMs computed in our study.
System SARI
Hybrid (Narayan and Gardent, 2014) $31.4$
DRESS-LS (Zhang and Lapata, 2017) $37.3$
PBSMT-R (Wubben et al., 2012) $38.6$
SBSMT-SARI (Xu et al., 2016) $40.0$
DMASS-DCSS (Zhao et al., 2018) $40.4$
Llama-3.2-3B $38.2$
Qwen2.5-72B $42.7$
GPT-4 $42.9$
The metrics’ scores are detailed in Table 9, with the best scores emphasized in bold.
SARI Shows Effectiveness on Transformation-Rich Datasets
SARI favors GPT-4 and Qwen2.5-72B over Llama-3.2-3B and Control-T5 in transformation-rich datasets — ASSET and SimPA, aligning with our human evaluations of overall superior performance. In Turk, which focuses on lexical paraphrasing, GPT-4, Qwen2.5-72B, and Control-T5 are scored similarly, all higher than Llama-3.2-3B. Considering that our error analysis points out that GPT-4 and Qwen2.5-72B tend to use more complex lexical expressions, SARI may preserve sensitivity on lexical simplicity, the same as what was reported in a previous study (Xu et al., 2016). In Newsela, GPT-4, Qwen2.5-72B and Llama-3.2-3B are scored similarly, all higher than Control-T5. Although overall GPT-4 and Qwen2.5-72B surpass Llama-3.2-3B in Newsela, our error analysis finds that sometimes Llama-3.2-3B merges multiple simplifications into a single output, resulting in repetitive content (see Section 6.1.3). This behavior may inflate Llama-3.2-3B’s SARI score on Newsela. Before the emergence of LLMs, sentence simplification learned transformation rules from parallel corpora of original-simplified sentence pairs (Alva-Manchego et al., 2020b). We compare LLM-generated simplifications with those from well-established models by examining SARI scores reported by Alva-Manchego et al. (Alva-Manchego et al., 2019) on the Turk Corpus. As shown in Table 9, larger LLMs such as Qwen2.5-72B and GPT-4 surpass all previously reported systems in terms of SARI scores, while the smaller Llama-3.2-3B model demonstrates comparable performance. These findings are consistent with our human evaluation results.
BLEU is Unsuitable
BLEU significantly favors Control-T5 on ASSET and Newsela, which does not match our human evaluations. Studies (Xu et al., 2016; Sulem et al., 2018a) have demonstrated that BLEU is unsuitable for simplification tasks, as it tends to negatively correlate with simplicity, often penalizing simpler sentences, and gives high scores to sentences that are close or even identical to the input. Our finding further underscores the limitations of BLEU in evaluating sentence simplification.
Limitations of FKGL in Comprehensive Quality Evaluation
@@
FKGL ranks Control-T5’s outputs as easier to read compared to those from GPT-4 and Qwen2.5-72B across all datasets. This aligns with our human evaluation; Control-T5 tends to generate simpler sentences at the expense of meaning preservation, thereby making the sentence easier to read. However, FKGL’s focus solely on readability, without taking into account the quality of the content or the reference sentences, limits its effectiveness in a comprehensive quality analysis. Previous studies (Alva-Manchego et al., 2021; Tanprasert and Kauchak, 2021) show that FKGL is unsuitable for sentence simplification evaluation. Our finding further highlights its limitations in accurately evaluating corpus-level sentence simplification.
@@ 7.4. Summary of Findings
We summarize our findings on the meta-evaluation of existing evaluation metrics for sentence simplification, namely LENS, BERTScore, SARI, BLEU, and FKGL.
@@
    At the overall sentence level, both LENS and BERTScore fail to effectively differentiate between high-quality and low-quality simplifications, particularly when evaluating high-quality simplification models, i.e., GPT-4 and Qwen2.5-72B.
@@
    For sentence-level meaning preservation, BERTScore is better at distinguishing high-quality from low-quality simplifications compared to LENS. Conversely, for sentence-level simplicity, LENS is more effective than BERTScore. However, the correlations of both metrics against human ratings are limited to the moderate range.
@@
    At the corpus level, SARI aligns more closely with our human evaluations, while BLEU appears less suitable.
@@ 8. Conclusion
In this study, we conduct an in-depth human evaluation of LLMs in sentence simplification. Our findings highlight that LLMs surpass the previous SOTA model, Control-T5, by generating fewer erroneous simplification outputs and preserving the source sentence’s meaning better. These results underscore the superiority of advanced LLMs in this task. Among LLMs, while larger LLMs generally perform better, performance does not always scale directly with size. Medium-sized LLMs may offer strong potential in simplification tasks. Nevertheless, we observed limitations in large and medium-sized LLMs, notably in GPT-4 and Qwen2.5-72B’s handling of lexical paraphrasing. Further, our meta-evaluation of sentence simplification’s automatic metrics demonstrates their inadequacy in accurately assessing the quality of LLM-generated simplifications.
With their advanced capabilities, LLMs hold great promise as tools for text simplification, benefiting non-native speakers and individuals with reading difficulties. However, the limitations identified in our study underscore the need for careful selection and application of these models to ensure they genuinely benefit users. For instance, models like Qwen2.5-72B and GPT-4 may be preferable to others. Even with these advanced models, caution is necessary to mitigate paraphrasing errors, as making lexical expressions more complex is counterproductive. We hope our study provides valuable insights for future research, contributing to improved simplification performance in LLMs and enhancing their usability and effectiveness across diverse user groups. Our investigation opens up multiple directions for future research. Future studies could investigate how to mitigate lexical paraphrasing issues. For example, it would be worthwhile to explore whether fine-tuning could help. Moreover, there’s a need for more sensitive automatic metrics to evaluate the sentence-level quality of simplifications generated by LLMs properly. Automating our error classification approach could enable real-time monitoring, allowing a plugin to periodically sample model-generated simplifications, identify error types and locations, and incorporate them into instructions to help the model self-correct.
@@ A.1. Best Prompts in GPT-4’s prompt engineering
Figure 9,  9, and  9 illustrate the best prompts that achieved the highest SARI scores on each validation set during GPT-4’s prompt engineering. Each prompt comprises: instructions, examples of original to simplification(s) transformation, and a source sentence.